# LCSim: A Large-Scale Controllable Traffic Simulator

### v1

@@ A.1 Problem Formulation
Similar to [23], we denote a traffic scenario as $\mathbf{\omega}=(M,A_{1:T})$, where $M$ contains the information of a High-Definition (HD) map and $A_{1:T}=[A_{1},...,A_{T}]$ is the state sequence of all traffic participates. Each element $m_{i}$ of $M=\{m^{1},...,m^{N_{m}}\}$ represents the map factor like road lines, road edges, centerline of lanes, etc. And each element $a_{i}^{t}$ of $A_{t}=\{a_{t}^{1},...,a_{t}^{N_{a}}\}$ represents the state of the ith traffic participate at time step t including position, velocity, heading, etc.
Given the map elements $M=\{m^{1},...,m^{N_{m}}\}$ and the historical states of agents $A_{t_{c}-T_{h}:t_{c}}$, where $T_{h}$ is the number historical steps and $0<T_{h}<t_{c}$, the model generates the future states of agents in the scenario $A_{t_{c}:t_{c}+T_{f}}$, where $T_{f}$ is the number of future steps.
A.2 Model Architecture
@@
Scene Encoder. We implemented our scene encoder based on MTR [32] and QCNet [51]. As mentioned before, at each time step $t_{c}$, the input to the scene encoder includes the map elements $M=\{m^{1},...,m^{N_{m}}\}$ and the historical states of agents $A_{t_{c}-T_{h}:t_{c}}$. First, we construct a heterogeneous graph $G=(V,E)$ based on the geometric relationships among input features. The node set $V$ contains two kinds of node $v^{a}$ and $v^{m}$ and the edge set $E$ consists of three kinds of edge $e^{a\to a}$, $e^{a\to m}$ and $e^{m\to m}$. Connectivity is established between nodes within a certain range of relative distances. For nodes like $v_{i}^{a}$ and $v_{j}^{m}$, their node features contain attributes independent of geographical location like lane type, agent type, agent velocity, etc. The position information of nodes is stored in the relative form within the edge features like $e_{ij}^{a\to m}=[\mathbf{p}_{j}^{m}-\mathbf{p}_{i}^{a},\mathbf{\theta}_{j}^{m}-\mathbf{\theta}_{i}^{a}]$, where $\mathbf{p}$ and $\mathbf{\theta}$ are position vector and heading angle of each node at current time step $t_{c}$. For each category of elements in the graph, we use an MLP to map their features into the latent space with dimension $N_{h}$ to get the node embedding $\mathbf{v}^{a}_{i,t}(t_{c}-T_{h}\leq t\leq t_{c})$, $\mathbf{v}_{j}^{m}$ and edge embedding $\mathbf{e}_{ij}^{a\to a}$, $\mathbf{e}_{ij}^{a\to m}$, $\mathbf{e}_{ij}^{m\to m}$. Then we apply four attention mechanisms in Table 4 to them to get the final scene embedding. The scene embedding consists of two components: the map embedding with a shape of $[M,N_{h}]$, and the agent embedding with a shape of $[A,T_{h},N_{h}]$.
Table 4: The attention mechanisms of scene encoder.
@@
Diffusion Decoder. Figure 8 shows the whole architecture of the diffusion decoder. Similar to [52], we implemented a DETR-like decoder to model the joint distribution of multi-agent action sequences. Denote the generation target as $\boldsymbol{x}\in\mathbb{R}^{A\times T_{f}\times N_{a}}$, which represents future $T_{f}$ steps’ actions of agents in the scenario. Firstly, noise $\boldsymbol{z}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\right)$ is added to the input sequence. Subsequently, the action sequence with noise for each agent is mapped to a latent space via an MLP, serving as the query embedding for that agent. The query is then added to the Fourier Embedding with noise level $\sigma$, similar to positional encoding, to inform the model about the current noise level. Next, the query vector undergoes cross-attention with map embeddings, embeddings of other agents in the scenario, and the historical state embedding of the current agent, resulting in a fused agent feature vector incorporating environmental information. Following this, self-attention is applied to the feature vectors of each agent to ensure the authenticity of interaction among the action sequences generated for each agent. Finally, the feature vectors from the latent space are mapped back to the agent’s action space via an MLP to obtain the de-noised agent action sequence.
A.3 Training Details
Table 5: Model parameters
Parameter Value
Input Size 2
Output Size 2
Embedding Size 128
Num Historical Steps 10
Num Future Steps 80
Num Polygon Types 20
Num Freq Bands 64
Map Encoder
   Hidden Dim 64
   Num Layers 5
   Num Pre Layers 3
Agent Encoder
   Time Span 10
   a2a Radius 50
   pl2a Radius 50
   Num Layers 2
   Num Heads 8
   Head Dim 64
   Dropout 0.1
Diff Decoder
   Output Head False
   Num t2m Steps 10
   pl2m Radius 150
   a2m Radius 150
   Num Layers 2
   Num Recurrent Steps 2
   Num Heads 8
   Head Dim 64
   Dropout 0.1
Table 6: Training parameters
Parameter Value
Batch Size 16
Num Epochs 200
Weight Decay 0.03
Learning Rate 0.0005
Learning Rate Schedule OneCycleLR
$\sigma_{data}$ 0.1
$c_{in}(\sigma)$ $1/\sqrt{\sigma^{2}+\sigma_{data}^{2}}$
$c_{skip}(\sigma)$ $\sigma_{data}^{2}/(\sigma^{2}+\sigma_{data}^{2})$
$c_{out}(\sigma)$ $\sigma\cdot\sigma_{data}/\sqrt{\sigma^{2}+\sigma_{data}^{2}}$
$c_{noise}(\sigma)$ $\frac{1}{4}\ln{\sigma}$
Noise Distribution $\ln(\sigma)\sim\mathcal{N}\left(P_{\text{mean }},P_{\text{std }}^{2}\right)$
$P_{\text{mean }}$1.2
$P_{\text{std }}$ 1.2
Training Target. Diffusion model estimates the distribution of generation target $\boldsymbol{x}\sim p(\boldsymbol{x})$ by sampling from $p_{\boldsymbol{\theta}}(\boldsymbol{x})$ with learnable model parameter $\boldsymbol{\theta}$. Normally we have $p_{\boldsymbol{\theta}}(\boldsymbol{x})=\frac{-f_{\boldsymbol{\theta}}(\boldsymbol{x})}{Z_{\boldsymbol{\theta}}}$ , and use max-likelihood $\max_{\boldsymbol{\theta}}\sum_{i=1}^{N}\log{p_{\boldsymbol{\theta}}(\boldsymbol{x}_{i})}$ to get parameter $\boldsymbol{\theta}$. However, to make the max likelihood training feasible, we need to know the normalization constant $Z_{\boldsymbol{\theta}}$, and either computing or approximating it would be a rather computationally expensive process, So we choose to model the score function $\nabla_{\boldsymbol{x}}\log{p_{\boldsymbol{\theta}}(\boldsymbol{x};\sigma)}$ rather than directly model the probability density, with the score function, one can get data sample $\boldsymbol{x}_{0}\sim p_{\boldsymbol{\theta}}(\boldsymbol{x})$ by the following equation [17]:
$\begin{array}[]{r}\boldsymbol{x}_{0}=\boldsymbol{x}(T)+\int_{T}^{0}-\dot{\sigma}(t)\sigma(t)\nabla_{\boldsymbol{x}}\log p_{\boldsymbol{\theta}}(\boldsymbol{x}(t);\sigma(t))dt\quad\text{ where }\boldsymbol{x}(T)\sim\mathcal{N}\left(\mathbf{0},\sigma_{\max}^{2}\boldsymbol{I}\right)\end{array}$ (1)
On this basis, we add a condition $\boldsymbol{c}$ composed of scene embeddings and use our model to approximate the score function $\nabla_{\boldsymbol{x}}\log p_{\boldsymbol{\theta}}(\boldsymbol{x};\boldsymbol{c},\sigma)\approx\left(D_{\boldsymbol{\theta}}(\boldsymbol{x};\boldsymbol{c},\sigma)-\boldsymbol{x}\right)/\sigma^{2}$, which leads to the training target [17]:
$\mathbb{E}_{\boldsymbol{x},\boldsymbol{c}\sim\chi_{c}}\mathbb{E}_{\sigma\sim q(\sigma)}\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\boldsymbol{I}\right)}\left\|D_{\boldsymbol{\theta}}(\boldsymbol{x}+\boldsymbol{\epsilon};\boldsymbol{c},\sigma)-\boldsymbol{x}\right\|_{2}^{2}$ (2)
$\chi_{c}$ is the training dataset combined with embeddings computed by the scene encoder, and $q(\sigma)$ represents the schedule of the noise level added to the original data sample. For better performance, we introduce the precondition as described in [18] to ensure that the input and output of the model both follow a standard normal distribution with unit variance:
$D_{\boldsymbol{\theta}}(\boldsymbol{x};\boldsymbol{c},\sigma)=c_{\text{skip }}(\sigma)\boldsymbol{x}+c_{\text{out }}(\sigma)F_{\boldsymbol{\theta}}\left(c_{\text{in }}(\sigma)\boldsymbol{x};\boldsymbol{c},c_{\text{noise }}(\sigma)\right)$ (3)
Here, $F_{\boldsymbol{\theta}}(\cdot)$ represents the original output of the diffusion decoder. In the experiment, we used the magnitude and direction of vehicle speed as the target for generation.
Experiment Setting. We trained our diffusion model on the Waymo Open Motion Dataset (WOMD) [26]. Each traffic scenario in the dataset has a duration of 9 seconds. We used the map information and the historical state of the previous 1 second as input to the model and generated future vehicle action sequences for the next 8 seconds. The training was conducted on a server with 4 $\times$ Nvidia 4090 GPUs. We set the batch size for training to 16 and trained with the OneCycleLR learning rate schedule for 200 epochs. The entire training process lasted approximately 20 days. The detailed parameters of the model and the training process are shown in Table 6 and Table 6.
A.4 Guide Functions
Following [50, 17], we calculate the cost function $\mathcal{L}:\mathbb{R}^{A\times T_{f}\times N_{a}}\mapsto\mathbb{R}$ based on the intermediate results of the generation process and propagate gradients backward to guide the final generation outcome. In our experiments, the control objectives include the vehicle’s maximum acceleration, target velocity, time headway, and relative distance to the preceding car during car-following, and generating adversarial behavior by controlling nearby vehicles to approach the current vehicle. Denote vehicle $i$ at timestep $t$ has states $acc_{i,t},v_{i,t},x_{i,t},y_{i,t},heading_{i,t}$, and $dis_{t}(i,j)$ computes the relative distance between vehicle $i$ and vehicle $j$ at timestep $t$ when vehicle $i$ is followed by vehicle $j$ on the same lane. Table 7 shows the details of the cost functions.
Table 7: The cost functions used in the guided generation process.
@@ A.4 Guide Functions
Appendix B Simulation System
B.1 Scenario Generator
We defined a unified map and vehicle Origin-Destination (OD) format based on Protobuf [^4] https://github.com/tsinghua-fib-lab/LCSim/blob/main/lcsim/protos. Additionally, we have developed format conversion tools designed for the Waymo and Argoverse datasets, the conversion results can be seen in Figure 9.
Figure 9: Traffic scenarios from WOMD (blue box) and Argoverse (yellow box).
B.2 Policy Details
We implemented five different policies to support traffic simulation in various scenarios:
  • BicycleExpertPolicy: Based on the expert policy, we impose kinematic constraints on the vehicle’s behavior using a bicycle model to prevent excessive acceleration and steering. By default, we set max acceleration to 6.0 $m/s^{2}$ and max steering angle to 0.3 $rad$.
  • RL-based Policy: A PPO [31] agent trained based on our simulator, its observation space contains the scene embedding and the action sequence. The action space consists of acceleration and steering values. The training environment of this agent is the second one, enabling diffusive simulation with Waymo-style vehicle behavior.
For the IDM model in these policies, the default configuration is that $acc_{max}=5m/s^{2},thw=2.0s,v_{target}=20m/s$.
Appendix C SenseTime Driving Dataset
C.1 Dataset Overview
SenseTime driving dataset comprises about 426.26 hours of vehicle driving logs collected from vehicles based on SenseAuto [^5] https://www.senseauto.com/ in the Beijing Yizhuang area and the whole dataset is split into 765 scenarios. The data is presented in a format similar to vehicle trajectories in the Waymo dataset with a sampled rate of 10 Hz. However, the road networks of the scenarios are not provided in this dataset, so we can not train our model on it, but due to the sufficient duration of the data, we can analyze the behavioral characteristics of vehicles within the data collection area. This analysis provides a reference for constructing driving scenarios with different styles.
Understandably, due to confidentiality regulations, the complete dataset cannot be released. However, we will share the statistical distribution data of vehicle behaviors obtained from the dataset.
C.2 Vehicle Behavior Analysis

### v2

@@ III-C Diffusion-based Motion Planner
We denote a traffic scenario as $\mathbf{\omega}=(M,A_{1:T})$ [11], where $M$ contains the information of a High-Definition (HD) map and $A_{1:T}=[A_{1},...,A_{T}]$ is the state sequence of all traffic participates. Each element $m_{i}$ of $M=\{m^{1},...,m^{N_{m}}\}$ represents the map factor like road lines, road edges, centerline of lanes, etc. And each element $a_{i}^{t}$ of $A_{t}=\{a_{t}^{1},...,a_{t}^{N_{a}}\}$ represents the state of the ith traffic participate at time step t including position, velocity, heading, etc. Given the map elements $M=\{m^{1},...,m^{N_{m}}\}$ and the historical states of agents $A_{t_{c}-T_{h}:t_{c}}$, where $T_{h}$ is the number historical steps and $0<T_{h}<t_{c}$, the model generates the future motion plans of vehicles $\bm{\tau}$ and computes future states of agents in the scenario $A_{t_{c}:t_{c}+T_{f}}$ from $\bm{\tau}$, where $T_{f}$ is the number of future steps. The entire model consists of the following three main components:
Algorithm 1 Generate Controllable Motion Plans
1:  Require diffusion decoder $D_{\bm{\theta}}$, scene embeddings $\bm{c}$, guide function $\mathcal{G}$, diffusion steps $t_{i\in\{0,...,N\}}$, guide gradient descent steps $K$, guide scale $\alpha$, guide clip $\beta$, initial noise level $S_{noise}$
2:  Initialize white noise $\bm{\tau}^{0}\sim\mathcal{N}(\bm{0},S_{noise}^{2}\bm{I})$
3:  for $i=0,\ldots,N$ do
4:     Denoising Step
5:     $\bm{d}_{i}=(\bm{\tau}^{i}-D_{\bm{\theta}}(\bm{\tau}^{i};\bm{c},t_{i}))/t_{i}$
6:     $\bm{\tau}^{i+1}=\bm{\tau}^{i}+(t_{i+1}-t_{i})\bm{d}_{i}$
7:     if $t_{i+1}\neq 0$ then
8:        $\bm{d}_{i}^{\prime}=(\bm{\tau}^{i+1}-D_{\bm{\theta}}(\bm{\tau}^{i+1};\bm{c},t_{i+1}))/t_{i+1}$
9:        $\bm{\tau}^{i+1}=\bm{\tau}^{i}+(t_{i+1}-t_{i})(\frac{1}{2}\bm{d}_{i}+\frac{1}{2}\bm{d}_{i}^{\prime})$
10:     end if
11:     Guide Step
12:     $\bm{\tau}^{i+1}_{0}=\bm{\tau}^{i+1}$
13:     for $j=1,\ldots,K$ do
14:        $\bm{\tau}^{i+1}_{j}=\bm{\tau}^{i+1}_{j-1}+\alpha\nabla\mathcal{G}(\bm{\tau}^{i+1}_{j-1})$
15:        $\Delta\bm{\tau}=|\bm{\tau}^{i+1}_{j}-\bm{\tau}^{i+1}_{0}|;\Delta\bm{\tau}\leftarrow clip(\Delta\bm{\tau},-\beta,\beta)$
16:        $\bm{\tau}^{i+1}_{j}\leftarrow\bm{\tau}^{i+1}_{0}+\Delta\bm{\tau}$
17:     end for
18:  end for
19:  Execute motion plans of each vehicle using output $\bm{\tau}^{N}_{K}$
III-C1 Scene Encoder
@@
We implemented our scene encoder based on MTR [50] and QCNet [51]. As mentioned before, at each time step $t_{c}$, the input to the scene encoder includes the map elements $M=\{m^{1},...,m^{N_{m}}\}$ and the historical states of agents $A_{t_{c}-T_{h}:t_{c}}$. First, we construct a heterogeneous graph $G=(V,E)$ based on the geometric relationships among input features. The node set $V$ contains two kinds of node $v^{a}$ and $v^{m}$ and the edge set $E$ consists of three kinds of edge $e^{a\to a}$, $e^{a\to m}$ and $e^{m\to m}$. Connectivity is established between nodes within a certain range of relative distances. For nodes like $v_{i}^{a}$ and $v_{j}^{m}$, their node features contain attributes independent of geographical location like lane type, agent type, agent velocity, etc. The position information of nodes is stored in the relative form within the edge features like $e_{ij}^{a\to m}=[\mathbf{p}_{j}^{m}-\mathbf{p}_{i}^{a},\mathbf{\theta}_{j}^{m}-\mathbf{\theta}_{i}^{a}]$, where $\mathbf{p}$ and $\mathbf{\theta}$ are position vector and heading angle of each node at current time step $t_{c}$. For each category of elements in the graph, we use an MLP to map their features into the latent space with dimension $N_{h}$ to get the node embedding $\mathbf{v}^{a}_{i,t}(t_{c}-T_{h}\leq t\leq t_{c})$, $\mathbf{v}_{j}^{m}$ and edge embedding $\mathbf{e}_{ij}^{a\to a}$, $\mathbf{e}_{ij}^{a\to m}$, $\mathbf{e}_{ij}^{m\to m}$. Then we apply four attention mechanisms in Table II to them to get the final scene embedding. The scene embedding consists of two components: the map embedding with a shape of $[M,N_{h}]$, and the agent embedding with a shape of $[A,T_{h},N_{h}]$.
TABLE II: The attention mechanisms of scene encoder.
@@ III-C2 Denoising Process
Figure 4 shows the whole architecture of the diffusion decoder. The vehicle’s future velocities and heading angles are used as the generation target for the diffusion model. Denote the generation target as $\bm{\tau}\in\mathbb{R}^{A\times T_{f}\times N_{a}}$, which represents future $T_{f}$ steps’ motion plans of vehicles in the scenario. Firstly, noise $\bm{z}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\right)$ is added to the input sequence. Subsequently, the noise motion plan for each vehicle is mapped to a latent space via an MLP, serving as the query embedding for that vehicle. The query is then added to the Fourier Embedding with noise level $\sigma$, similar to positional encoding, to inform the model about the current noise level. Next, the query vector undergoes cross-attention with map embeddings, embeddings of other vehicles in the scenario, and the historical state embedding of the current vehicle, resulting in a fused vehicle feature vector incorporating environmental information. Following this, self-attention is applied to the feature vectors of each vehicle to ensure the authenticity of interaction among the action sequences generated for each vehicle. Finally, the feature vectors from the latent space are mapped back to the vehicle’s action space via an MLP to obtain the denoised motion plans.
TABLE III: The cost functions used in the guided generation process.
@@
III-C3 Guide Function
Similar to CTG [21], we impose a loss function on the intermediate results of the denoising process and backpropagate the gradients to guide the generation process of the diffusion model. In our experiments, the control objectives include the vehicle’s maximum acceleration, target velocity, time headway, and relative distance to the preceding car during car-following, and generating adversarial behavior by controlling nearby vehicles to approach the current vehicle. Denote vehicle $i$ at timestep $t$ has states $acc_{i,t},v_{i,t},x_{i,t},y_{i,t},heading_{i,t}$, and $dis_{t}(i,j)$ computes the relative distance between vehicle $i$ and vehicle $j$ at timestep $t$ when vehicle $i$ is followed by vehicle $j$ on the same lane. Table III shows the details of the cost functions.
III-C4 Training Process
Diffusion model estimates the distribution of generation target $\bm{x}\sim p(\bm{x})$ by sampling from $p_{\bm{\theta}}(\bm{x})$ with learnable model parameter $\bm{\theta}$. Normally we have $p_{\bm{\theta}}(\bm{x})=\frac{-f_{\bm{\theta}}(\bm{x})}{Z_{\bm{\theta}}}$ , and use max-likelihood $\max_{\bm{\theta}}\sum_{i=1}^{N}\log{p_{\bm{\theta}}(\bm{x}_{i})}$ to get parameter $\bm{\theta}$. However, to make the max likelihood training feasible, we need to know the normalization constant $Z_{\bm{\theta}}$, and either computing or approximating it would be a rather computationally expensive process, So we choose to model the score function $\nabla_{\bm{x}}\log{p_{\bm{\theta}}(\bm{x};\sigma)}$ rather than directly model the probability density, with the score function, one can get data sample $\bm{x}_{0}\sim p_{\bm{\theta}}(\bm{x})$ by the following equation [52]:
$\begin{array}[]{l}\bm{x}_{0}=\bm{x}(T)+\int_{T}^{0}-\dot{\sigma}(t)\sigma(t)\nabla_{\bm{x}}\log p_{\bm{\theta}}(\bm{x}(t);\sigma(t))dt,\vspace{0.5em}\\ \text{where }\bm{x}(T)\sim\mathcal{N}\left(\mathbf{0},\sigma_{\max}^{2}\bm{I}\right).\end{array}$ (1)
On this basis, we add a condition $\bm{c}$ composed of scene embeddings and use our model to approximate the score function $\nabla_{\bm{x}}\log p_{\bm{\theta}}(\bm{x};\bm{c},\sigma)\approx\left(D_{\bm{\theta}}(\bm{x};\bm{c},\sigma)-\bm{x}\right)/\sigma^{2}$, which leads to the training target [52]:
$\mathbb{E}_{\bm{x},\bm{c}\sim\chi_{c}}\mathbb{E}_{\sigma\sim q(\sigma)}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\bm{I}\right)}\left\|D_{\bm{\theta}}(\bm{x}+\bm{\epsilon};\bm{c},\sigma)-\bm{x}\right\|_{2}^{2},$ (2)
here $\chi_{c}$ is the training dataset combined with embeddings computed by the scene encoder, and $q(\sigma)$ represents the schedule of the noise level added to the original data sample. For better performance, we introduce the precondition as described in [53] to ensure that the input and output of the model both follow a standard normal distribution with unit variance:
$D_{\bm{\theta}}(\bm{x};\bm{c},\sigma)=c_{\text{skip }}(\sigma)\bm{x}+c_{\text{out }}(\sigma)F_{\bm{\theta}}\left(c_{\text{in }}(\sigma)\bm{x};\bm{c},c_{\text{noise }}(\sigma)\right),$ (3)
here $F_{\bm{\theta}}(\cdot)$ represents the original output of the diffusion decoder. In the experiment, we used the magnitude and direction of vehicle speed as the target for generation. Our diffusion model can generate vehicle motion plans for 8 seconds in the future, we employ a recurrent generation approach based on a specified time interval during the simulation.
IV Experiments
We conduct a series of experiments to validate LCSim’s functionality. Firstly, we demonstrate how the diffusion-based motion planner in LCSim is constructed and its ability to create realistic and controllable traffic scenarios. Next, we train reinforcement learning agents in simulation environments with different driving styles built by LCSim, showcasing the impact of various driving environments on agent performance. Lastly, through the accurate replication of city-level traffic scenarios, we highlight LCSim’s capability to construct large-scale traffic simulations.
IV-A Performance of Diffusion-based Motion Planner
IV-A1 Datasets
We train our diffusion model on WOMD [24], which contains 500+ hours of driving logs collected from seven different cities in the United States. The dataset is further divided into scenario segments of 20s and 9s. In this experiment, we utilize the 9s segments for training, using the initial 1s as the historical context, and let the model generate vehicle motions in the future 8s. The diffusion model is trained on the training set and evaluated for its performance on the validation set.
IV-A2 Metrics
The evaluation metrics of the model consist of two aspects: first, the motion plans generated for vehicles by the model should adhere to basic traffic rules. We calculate the probabilities of vehicle collisions and off-road incidents in the simulated scenarios and compare these with statistical values from real-world data. Second, we use distance-based metrics to assess the model’s ability to reconstruct real-world traffic scenarios. For each scenario in the validation set, we sample K times for a simulation duration of T and compute the average displacement error across time (ADE) and the final displacement error (FDE) at the last time step. In the experiment, we set K = 6, and T = 8s, and choose the best matching sample to calculate minADE and minFDE.
IV-A3 Baselines & Settings
To validate the effectiveness of our model, we compare it with two famous learning-based traffic simulation methods, TrafficSim [43] and SimNet [18]. SimNet simulates traffic based on behavior cloning, while TrafficSim generates motion plans for vehicles using C-VAE. Additionally, to verify the effectiveness of the realistic guidance, we compare the model’s performance with and without the "no collision" and "no off-road" guidance. To handle the impact of randomness in the generation process, we conducted repeated experiments and reported the mean and error bars of the results.Baselines & Settings. To validate the effectiveness of our model, we compare it with two famous learning-based traffic simulation methods, TrafficSim [43] and SimNet [18]. SimNet simulates traffic based on behavior cloning, while TrafficSim generates motion plans for vehicles using C-VAE. Additionally, to verify the effectiveness of the realistic guidance, we compare the model’s performance with and without the "no collision" and "no off-road" guidance. To handle the impact of randomness in the generation process, we conducted repeated experiments and reported the mean and error bars of the results.
Figure 5: The process of generating vehicle motion plans by diffusion model.
TABLE IV: Evaluation results of our diffusion-based motion planner.
Collision (%) Off-Road (%) minADE (m) minFDE (m)
TrafficSim 4.901 ($\pm$ 0.019) 2.034 ($\pm$ 0.021) 1.205 ($\pm$ 0.001) 3.267 ($\pm$ 0.027)
SimNet 5.011 ($\pm$ 0.013) 1.996 ($\pm$ 0.017) 1.201 ($\pm$ 0.001) 3.259 ($\pm$ 0.025)
Ours w/o guide 9.693 ($\pm$ 0.413) 2.901 ($\pm$ 0.019) 1.383 ($\pm$ 0.002) 2.869 ($\pm$ 0.005)
Ours 4.118 ($\pm$ 0.082) 1.521 ($\pm$ 0.110) 1.526 ($\pm$ 0.005) 3.077 ($\pm$ 0.034)
Figure 6: The comparison of vehicle behaviors between WOMD and our diffusion-based simulation.
IV-A4 Results
Table IV presents the quantitative results of the experiments. While our method lags behind the baseline in terms of accurately imitating vehicle trajectories in the dataset, it surpasses the baseline in adherence to traffic rules, realistic vehicle interactions, and the accuracy of long-term simulations. Additionally, the comparison between simulations with and without realistic guidance demonstrates the effectiveness of the realistic guide functions. Furthermore, Figure 6 shows a comparison of vehicle behavior distributions collected from the original dataset and our diffusion-based simulations, indicating that our model can learn the driving styles present in the dataset. Figure 5 shows the visualization of the diffusion denoising process.
IV-B Alignment of Vehicle Behavior Characteristics
IV-B1 Private Driving Dataset